OpenStudy (anonymous):

Find the slope and y-intercept of each line whose equation is given. -4x + y = 7. Please show all of your work.

OpenStudy (anonymous):

I would manipulate this into slope intercept form so that it is obvious. y=mx+b where m is the slope and b is the y-intercept. add 4x to both sides y=4x+7 then pick out m... which is 4... that is slope then pick out b... which is 7... that is your y intercept

OpenStudy (anonymous):

ok so, y=4x+4x=7+4x?

OpenStudy (anonymous):

-4x+4x becomes zero in the process of balancing the equation I also see it as moving it to the other side.

OpenStudy (anonymous):

This is so confusing...-4x+4x=0+-4x+7=3

OpenStudy (anonymous):

I'll try to slow down a bit... If we start with -4x+y=7 and I want to make it y=stuff then I add 4x to both sides this cancels out the -4x on the left, but leaves 4x on the right it looks like -4x+4x+y=7+4x the next step becomes y=7+4x addition is commutative, so y=4x+7

OpenStudy (anonymous):

Where does the y=mx+b come in?

OpenStudy (anonymous):

This is the weird stuff... it is a formula that you read information from. As long as your equation has the y on the left... a number in front of the x... and just a constant add or subtracted, then it is in slope intercept form of a line. The number in front of the x (coefficient) is always the slope, and the number added or subtracted on the end is always the y-intercept.

OpenStudy (anonymous):

ok well I know you have others to help so I'm gonna try not to keep you, it's just that I can't seem to get it.

OpenStudy (anonymous):

I'm just a guy trying to help... nothing official. If you want more help from me, it is absolutely not a problem. If just think that someone else can explain it better... it won't hurt my feelings.

OpenStudy (anonymous):

I take classes online so I really don't have anyone to help me and I appreciate all you have done.

OpenStudy (anonymous):

No problem. There is an objective to this kind of problem. First, use balanced equation ideas to make it into a special form. Second, read the information from the form to get the information that you need. This is a new mode of math for most people who are used to simply calculating an answer from some numbers. This is analytical, meaning that you shape it, then read it.

OpenStudy (anonymous):

Lines always take a certain shape when in equation form. The official shape looks like this: Ax+By=C where A,B, and C are just numbers and x, and y are the variables. It may be official, but it is not very useful. y=mx+b says more. It is easy to plot a graph from it.

OpenStudy (anonymous):

So the first important part is... can you do the basic algebra to move stuff around in an equation? Then, memorize forms and meanings so that you can interpret. It looks like the first part about moving stuff around in the equation is still really hard for you. Do you still have questions about this problem or do you have another that you want to work through to see what I mean?

OpenStudy (anonymous):

People find it funny when I tell them that algebra wasn't a requirement in school over 28 years ago so, this is all very new to me and unfortunately I'm at work and don't have a lot of time to spend on this right now...Thanks for the help sweetie.

OpenStudy (anonymous):

I understand. Keep chipping away at it and you'll get it. Good Luck!